You have $375,000 invested at 7.5% and now you want to retire. You want to withdraw $35,000 at the end of each year, beginning at the end of this year. How many years will it take to exhaust your fund and run it down to $0?
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

375,000=35000[1-(1.075)^-time period]/0.075

375,000=466666.667[1-(1.075)^-time period]

(375,000/466666.667)=[1-(1.075)^-time period]

1-(375,000/466666.667)=(1.075)^-time period

(1/1.075)^time period=0.196428572

Taking log on both sides;

time period*log (1/1.075)=log 0.196428572

Time period=log 0.196428572/log (1/1.075)

=22.50 years(Approx)